Q1) Firm XXX has $4.5 million in current assets and 3.7 million in current liabilities. The firm wants to increase its inventory, which will be financed by a short term note with the bank. What level of inventories can the firm carry without its current ratio falling below 1.1?

Q2) Firm XXX has $4.5 million in current assets and 3.7 million in current liabilities. The firm wants to increase its inventory, which will be financed by issuing a 3-year bond. What level of inventories can the firm carry without its current ratio falling below 1.5?
Answer #1

1) Current Ratio = Current Asset/Current Liabilities
1.1 = (4.5+Inventories)/(3.7+Notes Payable)
1.1 = (4.5+x)/(3.7+x)
4.07 + 1.1x = 4.5 + x
Inventory can be increase by  = 0.39 million

2) Current Ratio = Current Asset/Current Liabilities ( 3 year bond is not part of Current Liabilities)
1.5 = (4.5+Inventories)/(3.7)
1.5= (4.5+x)/(3.7)
5.55 = 4.5+x
x = 1.55 million

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well